Sphere's volume is 900in^3. Find the Surface Area to the nearest whole number.
show all work and calculator functions

Answer by Boreal(15235)

V=(4/3)*pi*r^3=900 in^3
675=pi*r^3, multiplying both sides by (3/4)
calculator input 675, divide by pi and raise the answer to the (1/3) power.
5.99 cm
check with 6 cm
(4/3)(216*pi=288*pi=904.78 cm^3, which is slightly greater, as it should be.